Ophthalmology Glaucoma is an eye disease that is manifested by high intraocular pressure (IOP). The distribution of IOP in the general population is approximately normal with mean =16mmHg and standard deviation 3 mmHg. a. If the normal range for IOP is considered to be between 12 and 20 mmHg, then what percentage of the general population would fall within this range? b. If the top 5% of IOP is considered to be in danger range, then what minimum value of danger range? c. What is the 85% percentile of IOP?


1
Expert's answer
2022-03-14T17:49:46-0400

solution

Population mean "\mu =16"

Standard deviation "\sigma=3"


(a) Normal range (between 12 and 20)

We calculate Z values for values below 12 and above 20


"Z=\dfrac{X-\mu}{\sigma}"


"Z_{12}=\dfrac{12-16}{3}=-1.333"


"Z_{20}=\dfrac{20-16}{3}=1.333"


From the normal distribution tables

"P(Z_{12})=0.09176"

"P(Z_{20})=0.90824"


Since we are calculating the probability for values above "Z_{20}" we subtract the value from 1.

"=1-0.90824=0.09176"


Percentage in the normal range "(12 \to 20)"


"[1-(0.09176+0.09176)]\times100\%"

"=81.648\%"


(b)Top "5\%" of the population

"P=0.95"

From the normal distribution tables

"Z=1.65"

"Z=\dfrac{X-\mu}{\sigma}"


"X=Z\sigma+\mu"

"X=1.65\times3+16"

"X=20.95"


(c) 85th percentile of the population

"P=0.85"

From normal distribution tables

"Z=1.04"

"Z=\dfrac{X-\mu}{\sigma}"


"X=Z\sigma+\mu"

"X=1.04\times3+16"

"X=19.12"